# Exploring Large Language Models for Feature Selection: A Data-centric Perspective

### v1

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has significantly influenced various domains, leveraging their exceptional few-shot and zero-shot learning capabilities. In this work, we aim to explore and understand the LLMs-based feature selection methods from a data-centric perspective. We begin by categorizing existing feature selection methods with LLMs into two groups: data-driven feature selection which requires samples values to do statistical inference and text-based feature selection which utilizes prior knowledge of LLMs to do semantical associations using descriptive context. We conduct extensive experiments in both classification and regression tasks with LLMs in various sizes (e.g., GPT-4, ChatGPT and LLaMA-2). Our findings emphasize the effectiveness and robustness of text-based feature selection methods and showcase their potentials using a real-world medical application. We also discuss the challenges and future opportunities in employing LLMs for feature selection, offering insights for further research and development in this emerging field.
Large Language Models, Feature Selection
@@ 1 Introduction
Recent years have witnessed the remarkable development of Large Language Models (LLMs) (Achiam et al., 2023; Brown et al., 2020; Tan et al., 2024a; Touvron et al., 2023) across various domains and areas (Liang et al., 2022; Chang et al., 2023; Li et al., 2024c). By leveraging extensive training corpora and well-designed prompting strategies, LLMs demonstrate impressive few-shot and zero-shot capabilities in diverse tasks such as question answering (Wei et al., 2022; Wang et al., 2023b; Tong et al., 2024), information extraction (Wadhwa et al., 2023; Zhu et al., 2023) and knowledge discovery (Pan et al., 2024; Wang et al., 2024b, 2023a). The tuning-free nature also makes in-context learning (ICL) in LLMs achieve a great balance between efficiency and effectiveness (Tan et al., 2024b).
Feature selection (Dash & Liu, 1997; Li et al., 2017) is a critical data serving step that ensures relevant and high-quality data for downstream machine learning and data mining applications. While existing data-driven selection methods have achieved great success in scenarios with abundant data and metadata, there is an increasing demand for efficient feature selection methods with few or even zero samples for various reasons (Zhang et al., 2019). This need is particularly pronounced in sensitive applications such as predicting survival times for cancer patients (Tomczak et al., 2015; Wissel et al., 2022), where privacy concerns may prevent hospitals and patients from sharing their data, posing difficulties in the feature selection and engineering process. To address this challenge, recent studies (Jeong et al., 2024; Han et al., 2024) have explored leveraging the few-shot capability in LLMs to perform feature selection in low-resource settings and got promising results.
@@
In this work, our objective is to thoroughly explore and understand LLMs-based feature selection methods from a data-centric perspective. The conclusions and insights drawn from this exploration can provide insightful guidance for real-world applications where different types of resources and data are available. To begin with, we categorize the prompting strategies in previous studies (Choi et al., 2022; Jeong et al., 2024; Liu et al., 2024b; Han et al., 2024) for LLMs-based feature selection into two groups: (i) data-driven methods, which provide specific samples to LLMs (Liu et al., 2024b; Han et al., 2024), and (ii) text-based methods, which incorporate feature and task descriptions into the instruction (Choi et al., 2022; Jeong et al., 2024). These two prompting strategies require different data types: data-driven methods rely on sample points from datasets to do statistical inference while text-based methods need descriptive context for better semantic association between features and target variables. Figure 1 presents an overall comparison between the abovementioned methods and traditional feature selection algorithms. These differences make us curious about how LLMs perform with each of them under different data availability settings.
@@
Based on the abovementioned findings, we further explore the applicability of text-based feature selection with LLMs in a medical application. Specifically, we focus on the prediction of survival time for cancer patients (Tomczak et al., 2015; Wissel et al., 2022), which is a crucial task to evaluate both patient health and treatment effectiveness. To enhance the LLMs’ understanding of medical-specific gene names, we developed a Retrieval-Augmented Feature Selection (RAFS) method that leverages descriptions from the National Institutes of Health (NIH) as auxiliary context. Experiment results demonstrate our RAFS’s effectiveness in performing effective feature selection while safeguarding patient’s privacy. Finally, we outline the existing challenges and potential opportunities in employing LLMs for feature selection.
@@ 2.1 Feature Selection
Feature selection is the process of identifying and selecting the most relevant and important features or variables from a dataset to improve the performance and efficiency of a machine learning model (Dash & Liu, 1997; Guyon & Elisseeff, 2003; Chandrashekar & Sahin, 2014; Li et al., 2017). These feature selection methods can be generally categorized into three groups: filter, wrapper, and embedded approaches. Filter methods (Lazar et al., 2012) first rank features by performing correlation analysis and then selecting the most important ones for the following learning step. Typical filter methods include mutual information (Lewis, 1992; Ding & Peng, 2005), Fisher score (Hart et al., 2000; Gu et al., 2011) and maximum mean discrepancy (Song et al., 2012). By contrast, wrapper methods (Kohavi & John, 1997) use heuristic search strategies to identify a feature subset that optimally enhances the performance of certain prediction models (e.g., sequential selection (Luo & Chen, 2014) and recursive feature elimination  (Guyon et al., 2002)). For embedded approaches, it works together with specific machine learning models in the training phase by adding various regularization items in the loss function to encourage feature sparsity (Tibshirani, 1996; Yuan & Lin, 2006; Feng & Simon, 2017; Lemhadri et al., 2021).
@@ 2.2 Feature Selection with LLMs
There are already some works exploring the adaptation of LLMs in feature selection. (Choi et al., 2022) try to extract the relevant knowledge from LLMs as the task prior to performing feature selection, reinforcement learning and casual discovery. For feature selection, they design a prompt to instruct GPT-3 (Brown et al., 2020) to generate whether given features are important by answering “Yes” or “No”. Following them, (Jeong et al., 2024) expand the LLMs-based feature selection and propose three different pipelines that directly utilize the generated text output. They also conduct extensive experiments in evaluation across various model scales and prompting strategies. Besides, some studies devise more complex pipelines with LLMs in feature selection and feature engineering. (Liu et al., 2024b) introduce an In-Context Evolutionary Search (ICE-SEARCH) in Medical Predictive Analytics (MPA) applications. It involves recurrently optimizing the selected features by prompting LLMs to perform feature filtering based on test scores. (Han et al., 2024) employ LLMs as feature engineers to produce meta-features beyond the original features and combine them with simple machine learning models to improve predictions in downstream tasks. In this work, we aim to explore and understand LLMs in performing feature selection from a data perspective, offering further insights and hints for the adaptation of LLM-based feature selectors in real-world applications.
@@ 3 A Data-centric Taxonomy
Given a pre-trained LLM $M$, we follow the scoring-based method proposed by (Jeong et al., 2024), which prompt $M$ to generate an importance score $s_{i}$ for the given feature/ concept $f_{i}$ in the dataset $d$:
@@ 3.1 Data-driven Feature Selection
Recently, LLMs have been employed to directly handle numeric data, demonstrating their capabilities in numerical prediction and analytics (Gruver et al., 2024; Jin et al., 2023). Therefore, we build a data-driven feature selection method with LLMs by providing both features’ value $n_{f_{i}}$ and the value of the target variable $n_{y}$. Intuitively, LLMs are supposed to infer the correlation and perform statistical analysis to determine the importance of the given feature in the dataset.
To be more specific, assume there are $m$ samples available in the dataset $d$, we first build the sample pairs $SP_{i}$ using values of the chosen feature and target variable:
@@
Then, we curate the prompt $P_{f_{i}}$ using $SP_{i}$ and other instruction context $C$:
@@ 3.2 Text-based Feature Selection
Another line of work (Choi et al., 2022; Jeong et al., 2024) tries to employ the extensive semantics knowledge in LLMs (Li et al., 2024a) to perform feature selection. Specifically, they incorporate detailed dataset descriptions in the prompt, instructing LLMs to semantically distinguish the importance of a given feature using their inherent knowledge and experience.
@@ 4.1 Experiment Settings
  • LLaMA-2 (Touvron et al., 2023): 7B parameters.
  • LLaMA-2 (Touvron et al., 2023): 13B parameters.
  • ChatGPT (OpenAI, 2022): $\sim$175B parameters [^1] We use $\sim$ to denote the estimated size (Jeong et al., 2024) of the two closed-source LLMs.
  • GPT-4 (Achiam et al., 2023): $\sim$1.7T parametersfootnotemark: .
@@
  • Filtering by Mutual Information (MI) (Lewis, 1992).
  • Recursive Feature Elimination (RFE) (Guyon et al., 2002).
  • Minimum Redundancy Maximum Relevance selection (MRMR) (Ding & Peng, 2005).
@@
Figure 4: (a). Each feature selection method’s results in the classification task, categorized by different LLMs; for each method, we add an error bar to represent its standard variant among various data availabilities. (b). Each feature selection method’s results in the regression task, categorized by different LLMs. In each sub-figure, we include the average performance of traditional data-driven methods and the random selection method for comparison.
Datasets. In our evaluation, we consider both classification and regression tasks. For the classification task, we use six datasets: Adult (Asuncion et al., 2007), Bank (Moro et al., 2014), Communities (Redmond, 2009), Credit-g (Kadra et al., 2021), Heart [^2] https://kaggle.com/datasets/fedesoriano/heart-failure-prediction and Myocardial (Golovenkin et al., 2020). For the regression task, we use four datasets: Diabetes (Efron et al., 2004), NBA [^3] https://www.kaggle.com/datasets/bryanchungweather/nba-player-stats-dataset-for-the-2023-2024, Rideshare [^4] https://www.kaggle.com/datasets/aaronweymouth/nyc-rideshare-raw-data and Wine (Asuncion et al., 2007). Detailed statistics of datasets are given in Table 1.
Implementation Details. For each dataset, we fix the feature selection ratio to be 30%. We vary the data availability to conduct evaluations with 16-shot, 32-shot, 64-shot, and 128-shot configurations. The test performance is measured using a downstream L2-penalized logistic/linear regression model, selected via grid search with 5-fold cross-validation. We use the area under the ROC curve (AUROC) to evaluate classification tasks and mean absolute error (MAE) for regression tasks.
@@ 4.2 Result Analysis
Finding 2: Text-based feature selection with the most advanced LLMs can achieve comparable performance with traditional feature selection methods in every data availability setting. In Figure 3 (a), we observe that while GPT-4 with the text-based feature selection method performs slightly below the best traditional method (MRMR), it still demonstrates comparable performance, making it a competitive feature selection method in few-shot scenarios. However, when the LLM backbone is switched to smaller models, such as ChatGPT, there is a significant performance drop with the text-based selection method. Additionally, we conducted experiments on the full dataset using ’GPT-4 w/text’ alongside three traditional feature selection methods, and found that GPT-4 with the text-based method remains competitive even in the full-shot scenario.
Finding 3: Data-driven feature selection using LLMs struggles when number of samples increases. An interesting phenomenon we observed is a significant performance drop in the classification task when the sample size increases from 64 to 128 using the data-driven feature selection method (Figure 3 (b)). This drop is consistently observed across all four LLMs, indicating that each model generates poorer feature subsets as the sample size grows. We attribute this issue to LLMs struggling with processing long sequences, a challenge highlighted in many previous studies (Dong et al., 2023; Liu et al., 2024a). This limitation constrains the effectiveness of data-driven feature selection, which is why we did not include it in the full-shot experiment.
Finding 4: Text-based feature selection exhibits a stronger scaling law with model size compared to data-driven feature selection with LLMs. We investigated how scaling laws in model size affect feature selection capabilities. In Figure 3 (b), we observe a clear relationship between the size of LLMs and their text-based feature selection capabilities. In contrast, while GPT-4 shows significantly superior performance in data-driven feature selection, the other three LLMs do not clearly follow the scaling law. This suggests that text-based feature selection is a reliable approach that can be enhanced by using larger and more powerful LLMs.
@@ 5 Survival Prediction - A Case Study
We use a biomedical task to showcase the utilization of LLMs-based feature selection in real-world applications. Survival time prediction (Tomczak et al., 2015; Wissel et al., 2022) aims to predict cancer patients’ survival time based on their physical and physiological indicators, playing a critical role in patient risk management and boosting treatment selection. One of the significant challenges in survival prediction datasets is the huge volume of features (e.g., there are around 20,000 gene expression features in the TCGA (Tomczak et al., 2015) dataset). While previous studies performed data-driven feature selection methods such as principal component analysis (PCA) to address this issue (Wissel et al., 2023), as we mentioned in Section 1, It would cause serious privacy concerns for both patients and hospitals.
Impressed by the competitive performance and sample-free nature of text-based feature selection with LLMs, here we adopt it in the survival prediction application. In our preliminary experiments, we found LLMs have difficulties in directly understanding the domain-specific feature name (e.g., gene ID). Therefore, we borrow insights from retrieval-aug-mented generation (RAG) with LLMs (Gao et al., 2023; Chen et al., 2024; Li et al., 2024b) and propose Retrieval-Augmented Feature Selection (RAFS) to efficiently handle these biomedical-specific feature names. Specifically, we retrieve meta information about each feature name from the National Center for Biotechnology Information (NCBI) [^5] https://www.ncbi.nlm.nih.gov/ and provide this information to LLMs as the support document for a better feature selection.
@@ 5.1 Experiment Settings
We conduct experiments using the Lung Adenocarcinoma (LUAD) dataset in The Cancer Genome Atlas (TCGA) benchmark (Tomczak et al., 2015). Akin to (Wissel et al., 2023), we use clinical indicators and gene expression as the full feature set and fix the feature selection ratio to be 30%. We use PriorityLasso (Klau et al., 2018) as our machine learning backbone and report three metrics: Antolini’s Concordance (Antolini’s C) (Tomczak et al., 2015), Integrated Brier score (IBS) (Graf et al., 1999) and D-Calibration (D-CAL) (Haider et al., 2020), all of which are commonly-used metrics for survival prediction.

### v2

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has significantly influenced various domains, leveraging their exceptional few-shot and zero-shot learning capabilities. In this work, we aim to explore and understand the LLMs-based feature selection methods from a data-centric perspective. We begin by categorizing existing feature selection methods with LLMs into two groups: data-driven feature selection which requires numerical values of samples to do statistical inference and text-based feature selection which utilizes prior knowledge of LLMs to do semantical associations using descriptive context. We conduct experiments in both classification and regression tasks with LLMs in various sizes (e.g., GPT-4, ChatGPT and LLaMA-2). Our findings emphasize the effectiveness and robustness of text-based feature selection methods and showcase their potentials using a real-world medical application. We also discuss the challenges and future opportunities in employing LLMs for feature selection, offering insights for further research and development in this emerging field.
@@ 1 Introduction
Recent years have witnessed the remarkable development of Large Language Models (LLMs) [1, 4, 53, 58] across various domains and areas [37, 6, 36, 3]. By leveraging extensive training corpora and well-designed prompting strategies, LLMs demonstrate impressive few-shot and zero-shot capabilities in diverse tasks such as question answering [65, 64, 57], information extraction [60] and knowledge discovery [46, 63, 62]. The tuning-free nature also makes in-context learning (ICL) in LLMs achieve a great balance between efficiency and effectiveness [54].
Feature selection [10, 35] is a critical data serving step that ensures relevant and high-quality data for downstream machine learning and data mining applications. While existing data-driven selection methods have achieved great success in scenarios with abundant data and metadata, there is an increasing demand for efficient feature selection methods with few or even zero samples for various reasons [72]. This need is particularly pronounced in sensitive applications such as predicting survival times for cancer patients [56, 66], where privacy concerns may prevent hospitals and patients from sharing their data, posing difficulties in the feature selection and engineering process. To address this challenge, recent studies [26, 23] have explored leveraging the few-shot capability in LLMs to perform feature selection in low-resource settings and got promising results.
@@
In this work, our objective is to thoroughly explore and understand LLMs-based feature selection methods from a data-centric perspective. The conclusions and insights drawn from this exploration can provide insightful guidance for real-world applications where different types of resources and data are available. To begin with, we categorize the prompting strategies in previous studies [8, 26, 40, 23] for LLMs-based feature selection into two groups: (i) data-driven methods, which provide specific samples to LLMs [40, 23], and (ii) text-based methods, which incorporate feature and task descriptions into the instruction [8, 26]. These two prompting strategies require different data types: data-driven methods rely on sample points from datasets to do statistical inference while text-based methods need descriptive context for better semantic association between features and target variables. Figure 1 presents an overall comparison between the abovementioned methods and traditional feature selection algorithms. These differences make us curious about how LLMs perform with each of them under different data availability settings.
@@
Based on the abovementioned findings, we further explore the applicability of text-based feature selection with LLMs in a medical application. Specifically, we focus on the prediction of survival time for cancer patients [56, 66], which is a crucial task to evaluate both patient health and treatment effectiveness. To enhance the LLMs’ understanding of medical-specific gene names, we developed a Retrieval-Augmented Feature Selection (RAFS) method that leverages descriptions from the National Institutes of Health (NIH) as auxiliary context. Experiment results demonstrate our RAFS’s effectiveness in performing effective feature selection while safeguarding patient’s privacy. Finally, we outline the existing challenges and potential opportunities in employing LLMs for feature selection.
@@ 2.1 Feature Selection
Feature selection is the process of identifying and selecting the most relevant and important features or variables from a dataset to improve the performance and efficiency of a machine learning model [10, 20, 5, 35]. These feature selection methods can be generally categorized into three groups: filter, wrapper, and embedded approaches. Filter methods [31] first rank features by performing correlation analysis and then selecting the most important ones for the following learning step. Typical filter methods include mutual information [32, 11], Fisher score [24, 19] and maximum mean discrepancy [52]. By contrast, wrapper methods [30] use heuristic search strategies to identify a feature subset that optimally enhances the performance of certain prediction models (e.g., sequential selection [42] and recursive feature elimination  [21]). For embedded approaches, it works together with specific machine learning models in the training phase by adding various regularization items in the loss function to encourage feature sparsity [55, 71].
@@ 2.2 Feature Selection with LLMs
There are already some works exploring the adaptation of LLMs in feature selection. [8] try to extract the relevant knowledge from LLMs as the task prior to performing feature selection, reinforcement learning and casual discovery. For feature selection, they design a prompt to instruct GPT-3 [4] to generate whether given features are important by answering “Yes” or “No”. Following them, [26] expand the LLMs-based feature selection and propose three different pipelines that directly utilize the generated text output. They also conduct extensive experiments in evaluation across various model scales and prompting strategies. Besides, some studies devise more complex pipelines with LLMs in feature selection and feature engineering. [40] introduce an In-Context Evolutionary Search (ICE-SEARCH) in Medical Predictive Analytics (MPA) applications. It involves recurrently optimizing the selected features by prompting LLMs to perform feature filtering based on test scores. [23] employ LLMs as feature engineers to produce meta-features beyond the original features and combine them with simple machine learning models to improve predictions in downstream tasks. In this work, we aim to explore and understand LLMs in performing feature selection from a data perspective, offering further insights and hints for the adaptation of LLM-based feature selectors in real-world applications.
@@ 3 A Data-centric Taxonomy
Given a pre-trained LLM $M$, we follow the scoring-based method proposed by [26], which prompt $M$ to generate an importance score $s_{i}$ for the given feature/ concept $f_{i}$ in the dataset $d$:
@@ 3.1 Data-driven Feature Selection
Recently, LLMs have been employed to directly handle numeric data, demonstrating their capabilities in numerical prediction and analytics [18, 27]. Therefore, we build a data-driven feature selection method with LLMs by providing both features’ value $n_{f_{i}}$ and the value of the target variable $n_{y}$. Intuitively, LLMs are supposed to infer the correlation and perform statistical analysis to determine the importance of the given feature in the dataset.
To be more specific, assume there are $m$ samples available in the dataset $d$, we first build the sample pairs $SP_{i}$ using values of the $i_{th}$ feature and target variable:
@@
Then, we curate the prompt $P_{f_{i}}$ using $SP_{i}$ as few-shot examples and other instruction context $C$:
@@ 3.2 Text-based Feature Selection
Another line of work [8, 26] tries to employ the extensive semantics knowledge in LLMs [33] to perform feature selection. Specifically, they incorporate detailed dataset descriptions in the prompt, instructing LLMs to semantically distinguish the importance of a given feature using their inherent knowledge and experience.
@@ 4.1 Experiment Settings
  • LLaMA-2 [58]: 7B parameters.
  • LLaMA-2 [58]: 13B parameters.
  • ChatGPT [45]: $\sim$175B parameters111$\sim$ denotes the estimated size [26] of closed-source LLMs.
  • GPT-4 [1]: $\sim$1.7T parametersfootnotemark: .
@@
  • Filtering by Mutual Information (MI) [32].
  • Recursive Feature Elimination (RFE) [21].
  • Minimum Redundancy Maximum Relevance selection (MRMR) [11].
@@
Datasets. In our evaluation, we consider both classification and regression tasks. For the classification task, we use six datasets: Adult [2], Bank [44], Communities [50], Credit-g [28], Heart [^2] https://kaggle.com/datasets/fedesoriano/heart-failure-prediction and Myocardial [16]. For the regression task, we use four datasets: Diabetes [13], NBA [^3] https://www.kaggle.com/datasets/bryanchungweather/nba-player-stats-dataset-for-the-2023-2024, Rideshare [^4] https://www.kaggle.com/datasets/aaronweymouth/nyc-rideshare-raw-data and Wine [2]. Detailed statistics of datasets are given in Table 1.
Implementation Details. For each dataset, we fix the feature selection ratio to 30%. We vary the data availability for evaluations with 16-shot, 32-shot, 64-shot, and 128-shot configurations. The test performance is measured using a downstream L2-penalized logistic/ linear regression model, selected via grid search with 5-fold cross-validation. We use the area under the ROC curve (AUROC) to evaluate classification tasks and mean absolute error (MAE) for regression.
@@ 4.2 Result Analysis
Finding 2: Text-based feature selection with the most advanced LLMs achieves comparable performance with traditional feature selection methods in every data availability setting. In Figure 3 (a), we observe that while GPT-4 with the text-based feature selection method performs slightly below the best traditional method (MRMR), it still demonstrates comparable performance, making it a competitive feature selection method in few-shot scenarios. However, when the LLM backbone is switched to less capable models, such as LLaMA, the text-based selection method shows a significant performance drop. Additionally, we experiment on the full dataset using ‘GPT-4 w/text’ alongside three traditional feature selection methods, and found that GPT-4 with the text-based method remains competitive even in the full-shot scenario.
Figure 4: (a) Each feature selection method’s results in the classification task, categorized by different LLMs; for each method, we add an error bar to represent its standard variant among various data availabilities. (b) Each feature selection method’s results in the regression task, categorized by different LLMs. In each sub-figure, we include the average performance of traditional data-driven methods and the random selection method for comparison.
Finding 3: Data-driven feature selection using LLMs struggles when number of samples increases. An interesting phenomenon we observed is a significant performance drop in the classification task when the sample size increases from 64 to 128 using the data-driven feature selection method (Figure 3 (b)). This drop is consistently observed across all four LLMs, indicating that each model generates poorer feature subsets as the sample size grows. We attribute this issue to LLMs struggling with processing long sequences, a challenge highlighted in many previous studies [12, 39]. This limitation constrains the effectiveness of data-driven feature selection, which is why we did not include it in the full-shot experiment.
Finding 4: Text-based feature selection exhibits a stronger scaling law with model size compared to data-driven feature selection with LLMs. We investigated how scaling laws in model size affect feature selection capabilities. In Figure 3 (b), we observe a clear correlation between the size of LLMs and their text-based feature selection capabilities. In contrast, while GPT-4 shows significantly superior performance in data-driven feature selection, the other three LLMs do not clearly follow the scaling law. This suggests that text-based feature selection is a reliable approach that can be enhanced by using powerful LLMs.
@@ 5 Survival Prediction - A Case Study
We use a biomedical task to showcase the utilization of LLMs-based feature selection in real-world applications. Survival time prediction [56, 66] aims to predict cancer patients’ survival time based on their physical and physiological indicators, playing a critical role in patient risk management and boosting treatment selection. One of the significant challenges in survival prediction datasets is the huge volume of features (e.g., there are around 20,000 gene expression features in the TCGA [56] dataset). While previous studies performed data-driven feature selection methods such as principal component analysis (PCA) to address this issue [67], as we mentioned in Section 1, It would cause serious privacy concerns for both patients and hospitals.
Impressed by the competitive performance and sample-free nature of text-based feature selection with LLMs, here we adopt it in the survival prediction application. In our preliminary experiments, we found LLMs have difficulties in directly understanding the domain-specific feature name (e.g., gene ID). Therefore, we borrow insights from retrieval-aug-mented generation (RAG) with LLMs [15, 7, 34] and propose Retrieval-Augmented Feature Selection (RAFS) to efficiently handle these biomedical-specific feature names. Specifically, we retrieve meta information (e.g., official full name, summary and annotation information) about each feature name from the online National Center for Biotechnology Information (NCBI) [^5] https://www.ncbi.nlm.nih.gov/ and provide this information to LLMs as the support document for better feature selection.
@@ 5.1 Experiment Settings
We conduct experiments using the Lung Adenocarcinoma (LUAD) dataset in The Cancer Genome Atlas (TCGA) benchmark [56]. Akin to [67], we use clinical indicators and gene expression as the full feature set and fix the feature selection ratio to be 30%. We use PriorityLasso [29] as our machine learning backbone and report three metrics: Antolini’s Concordance (Antolini’s C) [56], Integrated Brier score (IBS) [17] and D-Calibration (D-CAL) [22], all of which are commonly-used metrics for survival prediction.